solve this vectors of OA=I+J+K ; AB=3I-2J+K; BC=I+2J-2K and CD =2I+J+3K FIND OD VECTOR

solve this     vectors of   OA=I+J+K ; AB=3I-2J+K; BC=I+2J-2K and  CD =2I+J+3K   
FIND OD  VECTOR
 

Grade:11

1 Answers

PASUPULETI GURU MAHESH
8 years ago
p { margin-bottom: 0.25cm; line-height: 120%; }

GIVEN VECTORS OA=I+J+K;AB=3I-2J+K;BC=I+2J-2K AND CD=2I+J+3K

WE KNOW THAT VECTOR AB =OB-OA

3I-2J+K= OB-( I+J+K)

OB=4I-J+2K

SIMILARLY BC =OC-OB

I+2J-2K=OC-(4I-J+2K)

OC=5I+J

 

CD=OD-OC

2I+J+3K = OD -(5I+J)

OD=7I+2J+3K

THERE FORE VECTOR OD 7I+2J+3K